OpenStudy (anonymous):

solve the seperable equation: dx/dt=3xt^2

OpenStudy (alexwee123):

integral?

OpenStudy (anonymous):

Differentiate..lol

OpenStudy (anonymous):

separate the variables (1/x)dx=3t^2dt integrate both sides can you do this ?

OpenStudy (anonymous):

i got it to ln(x)= t^3 +C

OpenStudy (unklerhaukus):

\[\frac{\text dx}{\text dt}=3xt^2\] \[\frac{\text dx}{x}=3t^2\text dt\] \[\int\frac{\text dx}{x}=\int3t^2\text dt\] \[\ln x=t^3+c\]

OpenStudy (anonymous):

how to i get from this to x=Ce^(t^3)

OpenStudy (unklerhaukus):

\[\ln x=t^3+c\]\[e^{\ln x}=e^{t^3+c}\]\[x=e^ce^{t^3}\]\[x=ke^{t^3}\]

OpenStudy (anonymous):

ok now you can write in explicit form raise both sides to e power \[\Large e^{\ln(x)}=e^{t^3+C}\] \[\Large x=e^ce^{t^3}\] or \[\Large x=Ce^{t^3}\]

OpenStudy (anonymous):

thanks